Calculate the kJ of heat per gram released by Pentaborane-9?

Pentaborane-9 (B5H9) is a colorless, highly-reactive liquid that will burst into flames when exposed to oxygen.The reaction is:

2B5H9 (l) + 12O2 (g) --> 5B2O3 (s) + 9H2O (l)

Calculate the kilojoules of heat released per gram of the compound reacted with oxygen.The standard enthalpy of formations of B5H9 (l), B2O3 (s), and H2O(l) are 42.84, -1271.94, and -285.83 kJ/mol, respectively.
Answer #1

ΔH = (5*-1271.94+9*-285.83-2*42.84)

ΔH = 9.017*103kJ/mole

so 12 moles of Oxygen is needed to get this energy,

for 1gm it will b = 9017/(12*32) = 23.48kJ/mole gm

answered by: CKLAIR
Answer #2
Step1 Heat of combustion =[ -1271.94x5 -9x285.83+2x42.84]/2

Step2 =[ -6359.7 -2572.47+85.68]/2 =-8846.49/2=-4423.245 kJ/mole
answered by: the game